53.A manufacturer of cereal finds that the masses of cereal in the companys 200-g packages are normally distributed with a mean of 200 g and a standard deviation of 16.3 g. a) What proportion of these boxes contain between 183.7 g and 216.3 g of cereal? b) What is the probability that a box selected at random contains more than 216.3 g of cereal? c) How many of a shipment of 120 box would you expect to contain between 178 g and 225 g of cereal? d) Determine the range of masses that you would expect 90% of these boxes of cereal to contain. Show steps
Answer #1

a)

µ =    200                              
σ =    16.30                              
we need to calculate probability for ,                                  
183.7   ≤ X ≤    216.3                          
X1 =    183.7   ,   X2 =   216.3                  
                                  
Z1 =   (X1 - µ ) / σ =   -1.000                          
Z2 =   (X2 - µ ) / σ =   1.000                          
                                  
P (   183.7   < X <    216.3   ) =    P (    -1   < Z <    1.000   )
                                  
= P ( Z <    1.000   ) - P ( Z <   -1.000   ) =    0.8413   -    0.1587   =    0.6827

b)

µ =    200                          
σ =    16.30                          
right tailed                              
X ≥   216.3                          
                              
Z =   (X - µ ) / σ =   1.00                      
                              
P(X ≥   216.3   ) = P(Z ≥   1.00   ) =   P ( Z <   -1.00   ) =    0.1587

c)

µ =    200                              
σ =    16.30                              
we need to calculate probability for ,                                  
178   ≤ X ≤    225                          
X1 =    178   ,   X2 =   225                  
                                  
Z1 =   (X1 - µ ) / σ =   -1.350                          
Z2 =   (X2 - µ ) / σ =   1.534                          
                                  
P (   178   < X <    225   ) =    P (    -1.349693252   < Z <    1.534   )
                                  
= P ( Z <    1.534   ) - P ( Z <   -1.350   ) =    0.9375   -    0.0886   =    0.8489

so, 0.8489 proportion of boxes contain between 178g and 225 g

so, total numbe r of boxes = 200*0.8489 = 169.779 (so, approximately 170 boxes)

d)

µ =    200                  
σ =    16.3                  
proportion=   0.90

so, we will calculate Z value for 0.05 , and 0.95
                      
Z value at    0.05   =   -1.645   (excel formula =NORMSINV(α))      
z=(x-µ)/σ                      
so, X=zσ+µ=   -1.645   *   16.3   +   200  
X   =   173.1889 g

Z value at    0.95   =   1.645   (excel formula =NORMSINV(α))      
z=(x-µ)/σ                      
so, X=zσ+µ=   1.645   *   16.3   +   200  
X   =   226.8111g

so, answer is (173.19g, 226.81g)